how do you solve this problem and what is the answer?
2(7n/4-1/6)=7/12

Answer by stanbon(75887)

2(7n/4-1/6)=7/12
-------
Divide both sides by 2 to get:
(7n/4 - 1/6) = 7/24
-------
Multiply thru by 24 to get:
6*7n - 4 = 7
----
42n = 11
n = 11/42
------------------
Cheers,
Stan H.
==================